# I Wish to Have an Argument: Argumentative Reasoning in Large Language Models

### v1

@@ 3.2 Performance and Number of Exemplars
On our second experiment we evaluated the performance of the LLMs in AM and APE, but compared with the number of exemplars seen by the LLMs. The number of exemplars was $\{0,4,8,16,\tau\}$, where $\tau$ is the maximum number allowed by the model’s token limit. In the case of GPT-3, $\tau$ tended to be around $4$; for GPT-4 it was a variable number depending on the task and setting, with an average of $44$ for symbolic, non-CoT approaches in AM ($15$ CoT, $9$ AMR) and $29$ concrete; and $22$ for both symbolic and concrete in APE. The total size of our collected data, including all exemplars, tasks, settings and five calls, is about 3.4GB. The results are in Figures 3 and 4. The models showed on average a peak in performance when the number of exemplars was $4$, and their performance steadily decreased as this number increased. This peak was independent of the task and the setting, since all curves were similar to one another. In the case of CoT, we did not observe such a trend.
Figure 3: Effect of the number of exemplars in AM. Thick blue line is the concrete approach, thickest red line is the average symbolic performance. Last entry in all plots is the average maximum number of exemplars supported. On average, the LLMs performance peaked at $4$ exemplars, and steadily decreased as the exemplars became more numerous. The exception to this observation was CoT prompting (solid red lines), where they consistently outperformed concrete prompting approaches and all but one of the symbolic approaches (dashed red lines), and did not show this decreasing trend. GPT-3’s performance was quickly outclassed by GPT-4, even under the same number of exemplars.
Figure 4: Effect of the number of exemplars in APE. Thick blue line is the concrete approach, thickest red line is the average symbolic performance. Last entry on each plot is the average maximum number of exemplars supported. Similar to AM, on average model performance peaked at $4$ exemplars, and steadily decreased as they became more numerous. We were unable to test CoT and AMR under this approach due to limitations on token length and budget. Dotted lines for symbolic are the various symbolic representations.
@@ 3.3 Performance and Input Representation
On our last experiment we explored whether the representation of the input text into progressively more abstract ways led to a change of performance in the symbolic performance of the models. In this section we focused solely on GPT-4 and AM, since the AMR graphs tended to be too big for the models to handle (for GPT-3 in both AM and APE, for GPT-4 only APE).
We tested three representations (text with indices inline, plain text, and with and without an AMR graph). There is no rigorous way to quantify the level of abstraction for these approaches. However, we consider the concrete approach to be least abstract, “hints” (indices inline, indices in output) to be slightly more abstract, and purely symbolic input representations (AMR graphs) as the most abstract. In between these categories, we ranked them based on the complexity of the output representation: BIO tags are more abstract than indices, since they have rules for matching, and both are more complex than concrete approaches since outputting the matching sentence is simpler than mapping to an arbitrary symbol and returning it. The full list of experiments is in Table 2, and our results in Figure 5.
In non-CoT approaches, we observed noticeable improvements when the abstraction level was sufficiently low (indices inline and indices in output; concrete), and as the abstraction increased, the model started to perform poorly, but retained above-random performance. These results are significant under a Welch’s $t$-test carried out on the predictions arrays: when the input text is unaltered, requesting line numbers or BIO tags as the output provides no difference in predictions with CoT ($p\approx 0.77$; large $p$-values imply the distributions have identical expected values), but does provide a noticeable difference without ($p<0.05$).
[…]
Task Input Output (label set)
AM* Text Text and |START|
AM* Text with indices Indices
AM* Text Indices
AM* Text BIO tags
AM Text with AMR graph BIO tags
AM AMR graph BIO tags
Table 2: Input representations tested, in roughly increasing order of abstraction. Tasks marked with an asterisk (*) were tested with and without CoT. The first row (input text, output text requested) corresponds to our concrete reasoning setting. Our ranking of abstraction is arbitrary: we consider the text with indices marked inline less abstract than a text without them, since the former provides a “hint” of what the label set is supposed to be like. Output representations with BIO tags as the output are more abstract, since it requires some rule matching in order to determine which symbol corresponds to which part of the problem. AMR graphs are considered the most abstract.
Figure 5: Effect of abstraction level in the input/output representation with respect to output performance in AM. For non-CoT approaches, the more abstract (under our criterion) the input representation, the more difficult it is for the model to solve the task. For example, if the input is conditioned to have “hints” about the desired output (for example, requesting indices and adding line numbers to the input text) the performance increases significantly. That said, even the most abstract representations (AMR graphs) outperform noticeably the random baseline. CoT approaches (at the top, marked with an asterisk) consistently outperformed their non-CoT counterparts, even when the maximum number of exemplars supported is much lower, and regardless of abstraction level. We did not test CoT with AMR in AM due to token limitations.
@@ 4 Discussion
4.1 Relationship of F-scores and Reasoning
The models, and GPT-4 in particular, performed well in both AM and APE, to the point of beating, or almost matching, the existing SOTA models. This alone is not sufficient to claim that the models are able to perform argumentative reasoning, let alone understand it.
When altering the input and output representation with conceptually minor changes–such as adding line numbers or requesting a BIO set as opposed to integers–the models had noticeably different performances. We argue that this is indicative of a potential false positive: the models are unable to reason in an argumentative setting, but give excellent appearance of being able to do so.
It is worth noting that, although GPT-4’s performance in APE is SOTA, it is still well-below the equivalent performance in AM. Although the tasks are slightly different, another plausible explanation is that the models failed to generalize due to the length of the task, which it is known to cause problems in transformer-based models (Anil et al., 2022).
4.2 Impact of CoT in Argumentative Reasoning
That said, we noticed that this performance difference did not extend to the inclusion of CoT. These prompts had on average better performance than their non-CoT counterparts, and we observed that it yielded better results in ill-posted (here, overly abstract) problems. However, our analysis showed that the output distributions for all CoT approaches were rather similar. This is perhaps indicative that CoT allows the models to return the desired output regardless of representation.
One possible explanation for the similar output distributions is due to the templatized nature of CoT: it usually starts with a common phrase (in our case, “Let’s read line-by-line and solve this step by step”) and follows very specific steps that represent (in our prompt) the reasoning required. While we don’t discard this possible explanation, it is worth noting that the input and output representations are completely different in all approaches. That said, it is clear that CoT mitigates the effect of input and output representations on the execution of the task.
4.3 The Exemplar Effect
Perhaps the most surprising aspect of our evaluation was the effect of the number of exemplars with respect to the model’s downstream performance. It was not a linear relationship, and instead there was a clear “sweet spot” at around $4$ samples. Originally, Brown et al. (2020) had indicated that more exemplars in GPT-3 would yield better performance, but that this number must be finetuned ahead of time. We observed the same thing, but these results did not hold for GPT-4. What this suggests, from our experiment results, is that (assuming that more exemplars imply better understanding of the task) the models are not accurately performing inference from the data provided.
That said, this “exemplar effect” did not extend to CoT settings, since the maximum number of tokens used had slightly higher performance than the other approaches. However, we do not discard the possibility that models with longer token limitations could also show a similar exemplar effect in CoT.
@@ 5.1 Reasoning Evaluation
It could be argued that our results are not complete in terms of evaluating argumentative reasoning capabilities. We agree: recognizing an argument is not the same as deciding whether it is sound, valid, or neither. However, without the ability of the model to show that it is able to recognize arguments and identify relations between them, any potentially generated argument or result evaluating the model’s performance in these tasks could be deemed as untrustworthy. Note that issues arising from the model not understanding the prompt (Webson and Pavlick, 2022) are not present in our evaluation, since the results reported here all perform better than the random baselines, and our analysis was comparative in nature.
5.2 Data contamination
@@
We factored out, to an extent, potential data contamination, which is known to impact downstream model performance (Carlini et al., 2023; Liang et al., 2022; Lee et al., 2023; De Wynter et al., 2023), by tasking the model to recognize arguments from the passage. However, we are unable to guarantee that the models have not been trained with this data, and therefore have at least some bias towards these results.
5.3 Model Setup
We only evaluated two models, so our results may not extend to other LLMs. We did not fine-tune the models, and opted instead to treating them as generalists performing in-context learning, in line with their standard, contemporary usage. However, it is likely that fine-tuning or specialized pre-training could show more insights into our results, and probably a more uniform performance distribution across representations: after all, MRC-APE and MLMC are specialist models and remain the SOTA in AM for RRv2.
6 Related Work
6.1 Informal Reasoning and Argumentation
In this section we limit ourselves to discuss informal reasoning and argumentation evaluation of LLMs. For an introduction to AM covering other modelling approaches see the survey by Lawrence and Reed (2020) and for a broader introduction to the field and open problems see Lauscher et al. (2022).
There has been work on using LLMs for AM (Zhao et al., 2023), as well as determining argument quality (van der Meer et al., 2022; Holtermann, Lauscher, and Ponzetto, 2022). More recently, Han et al. (2022) showed that GPT-3 is unable to mimic human-like inductive reasoning. While non-LLM-based approaches to APE do exist (Cheng et al., 2020; Bao et al., 2022), our work is the first to evaluate LLMs for this particular task.
A subset of informal reasoning, commonsense reasoning, is a very active of research. It is not strictly the same as argumentative reasoning, but it allows for the evaluation of some types of defeasible arguments. While not the direct subject of this work, there are reports that GPT-3’s performance is rather low in commonsense reasoning (Cong, 2022) and figurative language (Chakrabarty, Choi, and Shwartz, 2022). In addition, there are techniques to improve the model’s performance in this area, such as generating explanations for out-of-distribution observations (Brahman et al., 2021), corpora and benchmarks (Forbes et al., 2020; Rudinger et al., 2020; Pyatkin et al., 2023), and decoding (Cohen and Mooney, 2023) and prompting (Liu et al., 2022; Jin et al., 2022; Štefánik and Kadlčík, 2023; Lu et al., 2022) strategies.
Of note, Webson and Pavlick (2022) explored whether GPT-3 and other LLMs were able to understand the prompts. They found evidence to the contrary, and although they provided two alternate interpretations for their findings (the task was too hard or the model refused to answer), their main conclusion is that the models do not always follow, or understand, the instructions the way a human being would.
6.2 Formal Reasoning
The scientific community has long been interested in the evaluation of reasoning capabilities for statistical models, and LLMs are no exception ((Zhang et al., 2023); (Dziri et al., 2023); (Dan, Bastani, and Roth, 2022); (Liu et al., 2022); to name a few). Benchmarks like BIG-Bench Hard (Suzgun et al., 2023) have subsets of data that are meant to evaluate formal reasoning. Saparov and He (2023) performed a formal analysis of the reasoning capabilities of LLMs. They built an ontology and tested whether GPT-3 was able to perform modus ponens given a specific premise. They found that the models are unable to do so consistently, with a much lower performance under fictional worlds. This has been taken beyond formal logic and into real-world reasoning problems (e.g., maths problems), applying GPT-3, and other LLMs with much success (Open AI, 2023; Lewkowycz et al., 2022; Zong and Krishnamachari, 2023). That said, there exists evidence to the contrary on some tasks, such as arithmetic (Dziri et al., 2023). The latter reference also indicates that there is strong evidence that tasks with significant reasoning depths will cause these models to fail, a result also shared by Anil et al. (2022).
@@ 7 Conclusion
In this paper we evaluated the argumentative reasoning capabilities of GPT-3 and GPT-4. We altered the input and output representations of the data to try to understand their behavior via their performances in AM and APE. We found that the models present good performance in these tasks, but this was not sufficient to conclude that they were indeed performing any type of reasoning. On the contrary, our experiments showed that the input and output representations had significant impact on their downstream performance. This sensitivity suggested that the model’s applicability to critical uses such as data labelling and paper reviewing needs to be exercised with care.
However, when the input and output representations included CoT, our results became more surprising: the “exemplar effect” (too many exemplars being detrimental for performance) was no longer present, and the output distributions were very similar regardless of input and output representations. We were unable to conclude that CoT helped argumentative reasoning in LLMs, but we did note that it helped mitigate potential issues associated with ill-conditioned problems.
As mentioned in the Limitations section, our work may be expanded by exploring the ability of the models to working in-depth with argumentation. Our work did not evaluate their ability to judge an argument’s strength, or for the model to provide reasonable rebuttals. Moreover, due to token length limitations, we were unable to evaluate AMR, the most abstract setting we tested, with CoT. We believe that the performance difference between non-CoT AMR and CoT AMR would provide valuable insights on to what extent these models are able to discern abstract input representations successfully. All of these are well-known problems to the computational argumentation field and the scientific community at large. We plan to explore in further work these as they relate to LLMs (GPT-4 specifically, given its performance compared to GPT-3) as in-depth evaluation of these is important for downstream applications.
Finally, although the goal of our work is to study LLM argumentative reasoning capabilities, we believe that this type of representation-based analysis could be extended to other models and non-argument-related tasks, such as natural language inference.
@@ Acknowledgements
The authors wish to thank Liying Cheng for answering questions about the RRv2 dataset and its methodologies.
@@ Appendix A Full Prompts
 - I am not at all familiar with quantization methods
@@
  - If, however, the authors did a thorough job of
@@
 - We have updated our manuscript so that it states
@@
  - While our proposed DualPrecision is simple yet
@@
  - For the reviewers convenience, we re-iterate our
@@
  - Overall, we systematically analyze the model size

### v2

@@ 4.2 Performance and Number of Exemplars
We compared the number of exemplars ($\{0,4,8,16,\tau\}$, where $\tau$ is the maximum number) with the LLM performance. For GPT-3, $\tau$ tended to be around $4$; for GPT-4 it varied, with an average of $44$ for symbolic, non-CoT approaches in AM ($15$ CoT, $9$ AMR) and $29$ concrete; and $22$ for both approaches in APE. Results are in Figures 3(a) and 3(b). The LLMs peaked in performance at $4$ exemplars, and their scores decreased from there. This was independent of the task and setting. We did not observe this trend in CoT.
(a) Number of exemplars versus Fμ in AM. CoT (thin solid red lines) outperformed concrete and all but one of the symbolic approaches.
(b) Number of exemplars versus F01 in APE. We could not test CoT and AMR due to limitations on token length and budget.
Figure 3: Exemplar number versus score for AM and APE. In blue is the concrete approach; thick red line is the averaged performance (symbolic; dashed red lines). Last entry in all plots is the average maximum number of exemplars supported. LLM scoring in non-CoT peaked at around $4$ exemplars, and decreased afterwards.
@@ 4.3 Performance and Input Representation
In this section we focused on GPT-4 and AM and the following representations: text with indices inline, plain text, and with and without an AMR graph. There is no rigorous way to quantify the level of abstraction for these. However, we consider the concrete approach to be least abstract; “hints” (indices inline, indices in output) to be slightly more abstract; and purely symbolic input representations (AMR graphs) as most abstract. Other I/O representations are ranked based on the output representation: BIO tags are more abstract than indices (they require rules for matching); and both are more complex than concrete settings, since outputting a matching string is easier than mapping to an arbitrary symbol. The list of experiments is in Table 1, and our results in Figure 4.
For non-CoT approaches, we observed noticeable improvements in low-abstraction scenarios (indices inline and indices in output; concrete). As it rose, LLMs scored worse, though remaining above random. The results are significant under a Welch’s $t$-test on the prediction arrays. When the input text is unaltered, line numbers or BIO tags make no difference in predictions with CoT ($p\approx 0.77$; large $p$-values imply the distributions have identical expected values), but are noticeably worse without ($p<0.05$). Another $t$-test shows that $t<-0.86$ and $p\approx 0.39$ for inline indices when compared with its non-CoT version. Hence we reject the null hypothesis that the distributions are distinct. Since the performance in non-CoT was better on average (67% vs 70% Fμ), it is possible that CoT harms performance on outlier points. We compared CoT in other scenarios, and also observed large $p$ values when comparing with only indices in the output ($<0.98$); and when comparing the latter with BIO tags ($p<0.77$). This suggests that CoT is effective in highly abstract scenarios, such as requesting BIO tags, but detrimental otherwise.
Figure 4: Effect of abstraction in the I/O representation with respect to Fμ in AM. For non-CoT, the more abstract the input, the more difficult it is for the model to solve the task. Even the most abstract representations (AMR graphs) are noticeably better than a random. In CoT (top, marked with an asterisk) consistently outperformed non-CoT, even when the maximum number of exemplars supported is much lower, and regardless of abstraction level. We did not test CoT with AMR in AM due to token limitations.
@@ 5 Discussion
The LLMs scored highly in AM and APE, to the point of beating or almost matching the existing SOTA models. This is not sufficient to claim that the models are able to perform argumentative reasoning. When we altered the I/O representation with conceptually minor changes–such as adding line numbers or requesting a BIO set as opposed to integers–the LLMs had noticeably different performances. An explanation for the relatively low scores in APE is that the LLMs failed to generalize due to the length of the task; which causes problems in transformer-based models [21].
CoT prompts had on average higher scores than their non-CoT counterparts. They also yielded better scores in ill-posed (overly abstract) scenarios. However, our analysis showed that the output distributions for all CoT approaches were rather similar. This is perhaps indicative that CoT allows the models to return the same output regardless of representation. This appears to be due to the templatized nature of CoT, i.e., “Let’s read line-by-line and solve step-by-step”, and the specific steps needed to generate the output regardless of the I/O representation. This itself is not indicative of reasoning.
Finally, there is a clear peak at four exemplars with respect to the model’s downstream performance. This suggests that, assuming that more exemplars imply better information about the task, the LLMs are not accurately performing inference from the data provided. This exemplar effect did not extend to CoT settings, though we do not discard the possibility that models with longer token limitations could also show this trend. Overall, we pose that the models are unable to reason in an argumentative setting, but their scores give an excellent appearance of being able to do so.
@@ 6 Limitations
Our analysis has three main limitations. In terms of reasoning evaluation, it could be argued that our results are not complete in terms of evaluating argumentative reasoning capabilities. We agree: recognizing an argument is not the same as deciding its quality. However, without the ability of the model to show that it is able to recognize arguments and identify relations between them, any potentially generated argument or result evaluating the model’s performance in these tasks is untrustworthy.
@@
We factored out, to an extent, potential data contamination, which is known to impact downstream model performance [36, 37], by tasking the model to recognize arguments from the passage. However, we are unable to guarantee that the models have not been trained with this data, and therefore have at least some bias towards these results. Finally, we only evaluated two models, so our results may not extend to other LLMs. Likewise, we did not fine-tune the models, and opted instead to treating them as generalists performing in-context learning, in line with their contemporary usage.
@@ 7 Conclusion
We evaluated the argumentative reasoning capabilities of GPT-3 and GPT-4, by measuring whether they could recognize arguments from a passage–the first step on performing such reasoning. The LLMs score well in AM and APE, beating or nearly-matching the SOTA. However, statistical analysis on the LLMs’ predictions when subject to small, yet still human-readable, alterations in the I/O representations showed that the LLMs were extremely sensitive to the abstraction level and the number of exemplars. Hence, we concluded that they were not reasoning over the arguments seen.
However, symbolic prompting strategies (e.g. reducing the abstraction level of the prompt by adding line numbers) allowed the LLMs to score well and even beat CoT. We were also unable to conclude that CoT helped argumentative reasoning in LLMs, but did observe more robust results due to its templatized nature. Our analysis implies that it helps mitigate issues stemming from overly abstract or ill-conditioned problems.
As mentioned in Section 6, we did not evaluate the LLMs’ ability to judge an argument’s strength, or to provide reasonable rebuttals. Moreover, due to token length limitations, we were unable to evaluate AMR, the most abstract setting we tested, with CoT. We believe that this evaluation could provide valuable insights on to what extent these models are able to discern abstract input representations. Further work could explore these issues. Overall, our work suggests that LLM usage in areas like data labelling and paper reviewing must be exercised with care and good judgement.
@@ Acknowledgements
The authors wish to thank Liying Cheng for answering questions about the RRv2 dataset.
@@ Appendix A Full Prompts
- I am not at all familiar with quantization methods
@@
- If, however, the authors did a thorough job of
@@
- We have updated our manuscript so that it states
@@
- While our proposed DualPrecision is simple yet
@@
- For the reviewers convenience, we re-iterate our
@@
- Overall, we systematically analyze the model size